Consider the model of a market with four identical producers. Each producer has the same constant average variable cost. The aggregate demand for the product these firms produce is Y=1000.p-5. The producers choose their quantities simultaneously and the price is the highest price p at which the total quantity can be sold. The price at the equilibrium of this model is 20. What is the value of the firms’ constant average variable cost?
1
Expert's answer
2020-04-08T09:54:27-0400

"Y'=-5×\frac {1000}{p^6};"

"Y'=\frac {-5000}{p6};"

"Y'=\frac {-5000}{20^6};"

"-5000\u00d74=-20000;"